# Dissipative dynamics of an open quantum battery in the BTZ spacetime

### v1

@@ 3.2.1 $\Delta<A$ case
In Fig. 3 we show the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case ($\theta=0$ case). We fixed the energy-level spacing of the QB $\Delta\ell=0.1$, the charging amplitude $A\ell=0.2$, and the BTZ black hole mass $M=1$. Representative examples of different distance of the QB (or Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 3 (a), (b) and (c)), $r/r_{h}=1.2$ (Fig. 3 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 3 (g), (h) and (i)) have been chosen. The closed QB system driven by a static bias (see Eq. (30)) is also reported as a reference limit (see red curves). It is found that, in the presence of dissipation, the charged average energy presents a damped oscillatory behavior, whose amplitude is modulated by the exponential decay dictated by the incoherent relaxation and dephasing rates given in Eqs. (3.1). With the increase of the Hawking temperature, the damping strength is enhanced, and the average energy stored in the QB degrades faster to its asymptotic limit (3.1) for the Dirichlet boundary case, while it is opposite for the transparent and the Neumann boundary cases. The coupling strength $\mu\sqrt{\ell}$ could also enhance the dissipation as expected, showing stronger damping with the increase of the coupling strength. Besides, boundary conditions play an important role in the charging process. Although the boundary conditions do not change the average energy in the infinite charging time limit shown in (3.1), they could lead to different damping. We can see that the Neumann boundary condition causes stronger damping than the transparent one, while the transparent one induces stronger damping than the Dirichlet one. This difference is more clear when the Hawking temperature becomes smaller.
@@
For both the decoherence and the pure dephasing coupling cases, there is an optimal charging time, $\tau_{\text{opt}}$, at which the average energy stored in the QB takes the maxima. In Figs. 5 and 6 we plot this maximum average energy and the optimal charging time as a function of the Hawking temperature parameter $r/r_{h}$ by varying the energy-level spacing $\Delta\ell$ and the charging amplitude $A\ell$ for decoherence coupling case ($\theta=0$). We can see from Fig. 5 that the boundary condition influences the behavior of the maximum average energy and the optimal charging time vs the Hawking temperature parameter significantly. For the Dirichlet case, the maximum average energy and the optimal charging time increase with the decrease of Hawking temperature (Hawking temperature decreases with the increase of $r/r_{h}$, see Fig. 1). When the Hawking temperature approaches to zero, these two quantities approach to asymptotical value. However, we can find that for both the transparent case and the Neumann case, the maximum average energy and the optimal charging time decrease when the Hawking temperature decreases. These two quantities, when the Hawking temperature goes to zero, also approach to asymptotical value. Besides, no matter for which boundary conditions, we can find that when the charging amplitude $A\ell$ and the Hawking temperature are fixed, the maximum average energy increases with the increase of the energy-level spacing of the QB $\Delta\ell$, while the corresponding optimal charging time decreases. In Fig. 6 we fix the energy-level spacing of the QB, $\Delta\ell$, and plot the maximum average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle_{\text{max}}$ and the optimal charging time $\tau_{\text{opt}}$ as a function of the relative distance of the QB in BTZ spacetime $r/r_{h}$ for various charging amplitude $A\ell$. We can also find that the boundary condition influences the behavior of the maximum average energy and the optimal charging time vs the Hawking temperature parameter significantly. When the relative distance of the QB in BTZ spacetime $r/r_{h}$ approaches to the infinity, both the maximum average energy and the optimal charging time approach to asymptotical value. Furthermore, no matter for which boundary conditions, the maximum average energy increases with the increase of the charging amplitude, while the optimal charging time decreases when the energy-level spacing of the QB and the Hawking temperature are fixed.
@@
In Figs. 7 and 8 we plot the maximum average energy and the optimal charging time as a function of the Hawking temperature parameter $r/r_{h}$ by varying the energy-level spacing $\Delta\ell$ and the charging amplitude $A\ell$ for pure dephasing coupling case ($\theta=\pi/2$). We can find for this pure dephasing coupling case the maximum average energy and the optimal charging time share the same behaviors as that for the decoherence case shown in Figs. 5 and 6, while their magnitudes are different. This is because for both the decoherence case and the pure dephasing case their corresponding average energy associated to the QB share the same formula shown in (3.1) and (3.1), but with different incoherent relaxation and dephasing rates shown in (3.1) and (3.1).
@@ 3.2.2 $\Delta\geq\,A$ case
To consider the charging dynamics for $\Delta\geq\,A$ case, we in Fig. 9 plot the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case ($\theta=0$). We fix the energy-level spacing of the QB $\Delta\ell=0.13$, the charging amplitude $A\ell=0.1$, and the mass $M=1$. Representative examples of different distance of the QB (or Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 9 (a), (b) and (c)), $r/r_{h}=1.1$ (Fig. 9 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 9 (g), (h) and (i)) have been chosen. The closed QB system driven by a static bias (see Eq. (30)) is also reported as a reference limit (see red curves). It is found that, in the presence of dissipation, the charged average energy presents an oscillatory behavior in the early time, while it will finally approach a maximum for the long enough charging time if the Hawking temperature is high. This maximum average energy depends on the Hawking temperature, the charging amplitude $A\ell$, and the energy-level spacing of the QB $\Delta\ell$. Interestingly, unlike the $\Delta<A$ case, when $\Delta>A$ the maximum average energy might be more than that for the closed QB case if the Hawking temperature is high enough. Note that this advantage could occur for all the three different boundary condition cases, including the Dirichlet, the transparent, and the Neumann one. This means that the dissipation from vacuum field fluctuations in the BTZ spacetime might enhance the charging performance. In other words, through the interaction between the QB and the external quantum field one can in principle extract the energy from the vacuum fluctuations in curved space through dissipation in the charging process. Furthermore, we can also find that for the Dirichlet boundary case the average energy reaches a maximum earlier with the increase of the Hawking temperature. However, for the transparent boundary and the Neumann boundary cases when the Hawking temperature becomes higher, the average energy reaches the corresponding maximum later. Further comparing the average-energy dynamics for these three boundary condition cases, we can find that under the same conditions the average energy for the Neumann boundary case would be the first to reach a maximum value, then it is the transparent boundary case, and finally it is the Dirichlet one. We can also find that the stronger the coupling between the QB and the quantum field is, the earlier the QB reaches the maximum average energy during the charging process. However, the maximum value of the average energy does not depend on the coupling strength. This conclusion is valid for all the three boundary condition cases.
@@
In Fig. 10 we plot the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the pure dephasing coupling case ($\theta=\pi/2$). We also fixed the energy-level spacing of the QB $\Delta\ell=0.13$, the charging amplitude $A\ell=0.1$, and the mass $M=1$. Representative examples of different distance of the QB (or Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 10 (a), (b) and (c)), $r/r_{h}=1.1$ (Fig. 10 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 10 (g), (h) and (i)) have been chosen. We also choose the closed QB system driven by a static bias (see Eq. (30)) as a reference limit (see red curves). Since the average energy for the dephasing coupling case has the similar formula to that for the decoherence coupling case, seen from Eqs. (3.1) and (3.1), we can find that the average energy for the dephasing coupling case shares the similar dynamics to that for the decoherence coupling case discussed above. Specifically, although in the presence of dissipation the charged average energy presents an oscillatory behavior in the early time, finally it will reach a constant if the charging time is long enough. This constant average energy denotes the maximum average energy of the QB during the charging process if the Hawking radiation is high enough, satisfying $T\geq\Omega/\ln\big{[}\frac{\Delta\Omega+\Delta^{2}-A^{2}}{\Delta\Omega+A^{2}-\Delta^{2}}\big{]}$ as discussed above. It is interesting that the maximum average energy for this case is stored more in the QB than that for the closed QB case. This result is quite different compared with $\Delta<A$ scenario where the maximum average energy in the presence of dissipation would never be more than that for the closed QB case. For all the three boundary condition cases, this maximum average energy increases with the increase of the Hawking temperature, while it does not depend on the coupling strength between the QB and the external quantum field. However, the coupling strength would affect the time at which the average energy reaches the maximum. It is found that no matter for which boundary cases, the stronger the coupling is, the earlier the average energy reaches the maximum. Comparing the three different boundary cases, we can find for the Dirichlet boundary case the higher Hawking temperature may lead to the earlier maximum average energy that is reached by the QB, while this is quite the opposite for the other two boundary cases.
@@
We give a brief summary here: we can find that the dynamics of the average energy stored in the QB during the charging process depends on the quantum field boundary condition, the coupling form between the QB and external quantum field, Hawking temperature, the QB energy-level spacing, and the charging amplitude. It is shown that under certain conditions the maximum average energy stored in an open QB might be more than that for the closed QB case. This advantage could occur for all the three different boundary condition cases, including the Dirichlet, the transparent, and the Neumann one. It means that the dissipation from vacuum field fluctuations in the BTZ spacetime might enhance the charging performance.

### v4

@@ 3.2.1 $\Delta<A$ case
In Fig. 3 we show the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case ($\theta=0$ case). We fixed the energy-level spacing of the QB $\Delta\ell=0.1$, the charging amplitude $A\ell=0.2$, and the BTZ black hole mass $M=1$. Representative examples of different distance of the QB (or the local Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 3 (a), (b) and (c)), $r/r_{h}=1.2$ (Fig. 3 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 3 (g), (h) and (i)) have been chosen. The closed QB system driven by a static bias (see Eq. (30)) is also reported as a reference limit (see red curves). It is found that, in the presence of dissipation, the charged average energy presents a damped oscillatory behavior, whose amplitude is modulated by the exponential decay dictated by the incoherent relaxation and dephasing rates given in Eqs. (3.1). With the increase of the local Hawking temperature, the damping strength is enhanced, and the average energy stored in the QB degrades faster to its asymptotic limit (3.1) for the Dirichlet boundary case, while it is opposite for the transparent and the Neumann boundary cases. The coupling strength $\mu\sqrt{\ell}$ could also enhance the dissipation as expected, showing stronger damping with the increase of the coupling strength. Besides, boundary conditions play an important role in the charging process. Although the boundary conditions do not change the average energy in the infinite charging time limit shown in (3.1), they could lead to different damping. We can see that the Neumann boundary condition causes stronger damping than the transparent one, while the transparent one induces stronger damping than the Dirichlet one. This difference is more clear when the local Hawking temperature becomes smaller.
@@
For both the decoherence and the pure dephasing coupling cases, there is an optimal charging time, $\tau_{\text{opt}}$, at which the average energy stored in the QB takes the maxima. In Figs. 5 and 6 we plot this maximum average energy and the optimal charging time as a function of the local Hawking temperature parameter $r/r_{h}$ by varying the energy-level spacing $\Delta\ell$ and the charging amplitude $A\ell$ for decoherence coupling case ($\theta=0$). We can see from Fig. 5 that the boundary condition influences the behavior of the maximum average energy and the optimal charging time vs the local Hawking temperature parameter significantly. For the Dirichlet case, the maximum average energy and the optimal charging time increase with the decrease of the local Hawking temperature (local Hawking temperature decreases with the increase of $r/r_{h}$, see Fig. 1). When the local Hawking temperature approaches to zero, these two quantities approach to asymptotical value. However, we can find that for both the transparent case and the Neumann case, the maximum average energy and the optimal charging time decrease when the local Hawking temperature decreases. These two quantities, when the local Hawking temperature goes to zero, also approach to asymptotical value. Besides, no matter for which boundary conditions, we can find that when the charging amplitude $A\ell$ and the local Hawking temperature are fixed, the maximum average energy increases with the increase of the energy-level spacing of the QB $\Delta\ell$, while the corresponding optimal charging time decreases. In Fig. 6 we fix the energy-level spacing of the QB, $\Delta\ell$, and plot the maximum average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle_{\text{max}}$ and the optimal charging time $\tau_{\text{opt}}$ as a function of the relative distance of the QB in BTZ spacetime $r/r_{h}$ for various charging amplitude $A\ell$. We can also find that the boundary condition influences the behavior of the maximum average energy and the optimal charging time vs the local Hawking temperature parameter significantly. When the relative distance of the QB in BTZ spacetime $r/r_{h}$ approaches to the infinity, both the maximum average energy and the optimal charging time approach to asymptotical value. Furthermore, no matter for which boundary conditions, the maximum average energy increases with the increase of the charging amplitude, while the optimal charging time decreases when the energy-level spacing of the QB and the local Hawking temperature are fixed.
@@
In Figs. 7 and 8 we plot the maximum average energy and the optimal charging time as a function of the local Hawking temperature parameter $r/r_{h}$ by varying the energy-level spacing $\Delta\ell$ and the charging amplitude $A\ell$ for pure dephasing coupling case ($\theta=\pi/2$). We can find for this pure dephasing coupling case the maximum average energy and the optimal charging time share the same behaviors as that for the decoherence case shown in Figs. 5 and 6, while their magnitudes are different. This is because for both the decoherence case and the pure dephasing case their corresponding average energy associated to the QB share the same formula shown in (3.1) and (3.1), but with different incoherent relaxation and dephasing rates shown in (3.1) and (3.1).
@@ 3.2.2 $\Delta\geq\,A$ case
To consider the charging dynamics for $\Delta\geq\,A$ case, we in Fig. 9 plot the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case ($\theta=0$). We fix the energy-level spacing of the QB $\Delta\ell=0.13$, the charging amplitude $A\ell=0.1$, and the mass $M=1$. Representative examples of different distance of the QB (or local Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 9 (a), (b) and (c)), $r/r_{h}=1.1$ (Fig. 9 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 9 (g), (h) and (i)) have been chosen. The closed QB system driven by a static bias (see Eq. (30)) is also reported as a reference limit (see red curves). It is found that, in the presence of dissipation, the charged average energy presents an oscillatory behavior in the early time, while it will finally approach a maximum for the long enough charging time if the local Hawking temperature is high. This maximum average energy depends on the local Hawking temperature, the charging amplitude $A\ell$, and the energy-level spacing of the QB $\Delta\ell$. Interestingly, unlike the $\Delta<A$ case, when $\Delta>A$ the maximum average energy might be more than that for the closed QB case if the local Hawking temperature is high enough. Note that this advantage could occur for all the three different boundary condition cases, including the Dirichlet, the transparent, and the Neumann one. This means that the dissipation from vacuum field fluctuations in the BTZ spacetime might enhance the charging performance. In other words, through the interaction between the QB and the external quantum field one can in principle extract the energy from the vacuum fluctuations in curved space through dissipation in the charging process. Furthermore, we can also find that for the Dirichlet boundary case the average energy reaches a maximum earlier with the increase of the local Hawking temperature. However, for the transparent boundary and the Neumann boundary cases when the local Hawking temperature becomes higher, the average energy reaches the corresponding maximum later. Further comparing the average-energy dynamics for these three boundary condition cases, we can find that under the same conditions the average energy for the Neumann boundary case would be the first to reach a maximum value, then it is the transparent boundary case, and finally it is the Dirichlet one. We can also find that the stronger the coupling between the QB and the quantum field is, the earlier the QB reaches the maximum average energy during the charging process. However, the maximum value of the average energy does not depend on the coupling strength. This conclusion is valid for all the three boundary condition cases.
@@
In Fig. 10 we plot the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the pure dephasing coupling case ($\theta=\pi/2$). We also fixed the energy-level spacing of the QB $\Delta\ell=0.13$, the charging amplitude $A\ell=0.1$, and the mass $M=1$. Representative examples of different distance of the QB (or local Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 10 (a), (b) and (c)), $r/r_{h}=1.1$ (Fig. 10 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 10 (g), (h) and (i)) have been chosen. We also choose the closed QB system driven by a static bias (see Eq. (30)) as a reference limit (see red curves). Since the average energy for the dephasing coupling case has the similar formula to that for the decoherence coupling case, seen from Eqs. (3.1) and (3.1), we can find that the average energy for the dephasing coupling case shares the similar dynamics to that for the decoherence coupling case discussed above. Specifically, although in the presence of dissipation the charged average energy presents an oscillatory behavior in the early time, finally it will reach a constant if the charging time is long enough. This constant average energy denotes the maximum average energy of the QB during the charging process if the Hawking radiation is high enough, satisfying $T\geq\Omega/\ln\big{[}\frac{\Delta\Omega+\Delta^{2}-A^{2}}{\Delta\Omega+A^{2}-\Delta^{2}}\big{]}$ as discussed above. It is interesting that the maximum average energy for this case is stored more in the QB than that for the closed QB case. This result is quite different compared with $\Delta<A$ scenario where the maximum average energy in the presence of dissipation would never be more than that for the closed QB case. For all the three boundary condition cases, this maximum average energy increases with the increase of the local Hawking temperature, while it does not depend on the coupling strength between the QB and the external quantum field. However, the coupling strength would affect the time at which the average energy reaches the maximum. It is found that no matter for which boundary cases, the stronger the coupling is, the earlier the average energy reaches the maximum. Comparing the three different boundary cases, we can find for the Dirichlet boundary case the higher local Hawking temperature may lead to the earlier maximum average energy that is reached by the QB, while this is quite the opposite for the other two boundary cases.
@@
We give a brief summary here: we can find that the dynamics of the average energy stored in the QB during the charging process depends on the quantum field boundary condition, the coupling form between the QB and external quantum field, the local Hawking temperature, the QB energy-level spacing, and the charging amplitude. It is shown that under certain conditions the maximum average energy stored in an open QB might be more than that for the closed QB case. This advantage could occur for all the three different boundary condition cases, including the Dirichlet, the transparent, and the Neumann one. It means that the dissipation from vacuum field fluctuations in the BTZ spacetime might enhance the charging performance.
@@ 3.3 Direct comparison of $\langle\,E\rangle$ for different coupling form
Figure 12: Time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case with $\theta=0$. Three different boundary conditions in BTZ spacetime are shown for various mass of black hole $M$. Here we have taken the energy-level spacing $\Delta\ell=0.1$, the charging amplitude $A\ell=0.2$, and the coupling strength $\mu\sqrt{\ell}=0.1$. The first, second, and third row panels have taken the relative distance of the QB in BTZ spacetime $r/r_{h}=1.001$, $r/r_{h}=1.2$ and $r/r_{h}=1.8$, respectively.
3.4 Effect of mass of BTZ black hole on average energy
We have studied the dynamics of the average energy stored in an open QB that is located in the BTZ spacetime in the above sections. Where we fixed the mass of the black hole and mainly considered how the effective local black hole temperature, the driving amplitude, the boundary conditions on quantum field, and different dissipative couplings affect the relevant dynamics. Since small-mass black holes can produce interesting effects in relativistic quantum information, such as Fisher information FIBH , it is interesting to see what happens for the charging performances of a QB when the mass of black hole is small.
3.4.1 $\Delta<A$ case
Figure 13: Time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the pure dephasing coupling case with $\theta=\pi/2$. Three different boundary conditions in BTZ spacetime are shown for various mass of black hole $M$. Here we have taken the energy level spacing $\Delta\ell=0.1$, the charging amplitude $A\ell=0.2$, and the coupling strength $\mu\sqrt{\ell}=0.1$. The first, second, and third row panels have taken the relative distance of the QB in BTZ spacetime $r/r_{h}=1.001$, $r/r_{h}=1.2$ and $r/r_{h}=1.8$, respectively.
In Fig. 12 we show the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case ($\theta=0$ case) with different black hole masses $M=0.001$, $M=0.01$ and $M=1$. We fixed the energy-level spacing of the QB $\Delta\ell=0.1$, the charging amplitude $A\ell=0.2$, and the coupling strength $\mu\sqrt{\ell}=0.1$. Representative examples of different distance of the QB (or local Hawking temperature) away from the horizon $r/r_{h}=1.001$ (Fig. 12 (a), (b) and (c)), $r/r_{h}=1.2$ (Fig. 12 (d), (e) and (f)), and $r/r_{h}=1.8$ (Fig. 12 (g), (h) and (i)) have been chosen. It is found that, in the presence of dissipation, the charged average energy presents a damped oscillatory behavior, whose amplitude is modulated by the exponential decay dictated by the incoherent relaxation and dephasing rates given in Eqs. (3.1). When fixing the effective local Hawking temperature, we can find that with the increase of the black hole mass $M$, the damping strength is degraded. It means that the average energy stored in the QB degrades faster to its asymptotic limit (3.1) for smaller black hole mass. However, the smaller the black hole mass is, the less the maximum average energy stored in the QB during the charging process is. Therefore, if the local Hawking temperature is fixed, the mass of black hole can not lead to qualitatively different dynamical behavior of the average energy stored in the QB, this is different from the Fisher information case FIBH where small black hole mass can induce qualitative difference of the Fisher information dynamics. We can understand this from the expression of the average energy shown in Eq. (3.1). If the local Hawking temperature is fixed, the mass of black hole can only change the incoherent relaxation and dephasing rates given in Eqs. (3.1), thus can only change the damping strength. Besides, no matter for which black hole mass, the local Hawking temperature can enhance the damping of the dynamics of the average energy that is stored in the QB for the Dirichlet boundary case, while for the transparent and the Neumann boundary cases, the local Hawking temperature would degrade the damping of the dynamics of the average energy. Under the same condition, we can see that the Neumann boundary condition causes stronger damping than the transparent one, while the transparent one induces stronger damping than the Dirichlet one.
In Fig. 13 we show the time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the pure dephasing coupling case ($\theta=\pi/2$) with different black hole masses $M=0.001$, $M=0.01$ and $M=1$. As comparison, we fixed the same other parameters as that in Fig. 12. We find that in the presence of dissipation all curves present a damped oscillatory behaviour, whose amplitude is modulated by the exponential decay dictated by the incoherent relaxation and dephasing rates given in Eqs. (3.1). With the decrease of the black hole mass, the dynamics of the average energy presents a stronger damping. Which means that the smaller the black hole mass is, the faster the average energy approaches its asymptotical limit shown in (3.1). However, the smaller the black hole mass is, the smaller of the maximum average energy during the charging process is.
From the above analysis, we can find that the mass of black hole just changes the incoherent relaxation and dephasing rates shown in the Eqs. (3.1) and (3.1), thus only modifies damping strength of the dynamics of the average energy stored in the QB. Actually, one can arrive at this conclusion directly from the expressions of the average energy shown in (3.1) and (3.1).
Figure 14: Time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the decoherence coupling case ($\theta=0$). Three different boundary condition cases in BTZ spacetime are shown for various black hole mass $M=0.001$, $M=0.01$, and $M=1$. Here we have taken the energy level spacing $\Delta\ell=0.13$, the charging amplitude $A\ell=0.1$, and the coupling strength $\mu\sqrt{\ell}=0.1$. The first, second, and third row panels have taken the relative distance of the QB in BTZ spacetime $r/r_{h}=1.001$, $r/r_{h}=1.1$ and $r/r_{h}=1.8$, respectively.
Figure 15: Time evolution of the average energy stored in the QB $\langle\,E_{\text{QB}}(\tau)\rangle$ shown in Eq. (3.1) for the pure dephasing coupling case ($\theta=\pi/2$). Three different boundary condition cases in BTZ spacetime are shown for various black hole mass $M=0.001$, $M=0.01$, and $M=1$. Here we have taken the energy level spacing $\Delta\ell=0.13$, the charging amplitude $A\ell=0.1$, and the coupling strength $\mu\sqrt{\ell}=0.1$. The first, second, and third row panels have taken the relative distance of the QB in BTZ spacetime $r/r_{h}=1.001$, $r/r_{h}=1.1$ and $r/r_{h}=1.8$, respectively.
3.4.2 $\Delta\geq\,A$ case
In the above section we have analyzed how the black hole mass $M$ affects the average energy stored in the QB for the $\Delta<A$ case. Here we will discuss what happens for the average energy of QB when $\Delta\geq\,A$ in the following.
We fixed the coupled strength $\mu\sqrt{\ell}=0.1$, the energy level spacing $\Delta\ell=0.13$, effective local Hawking temperature, and the charging amplitude $A\ell=0.1$, and plotted the dynamics of the average energy of QB with different black hole mass $M=0.001$, $M=0.01$, and $M=1$ in Figs. 14 and 15, which denote the decoherence coupling case ($\theta=0$) and the pure dephasing coupling case ($\theta=\pi/2$), respectively. We can find that when we fix the other parameters, the mass of the black hole can not lead to qualitatively different dynamical behavior of the average energy. All the average energy presents oscillatory behavior with respect to the evolution time, and finally approaches to an asymptotical value at the infinite time limit. We can find that when the local Hawking temperature is fixed, this asymptotical value of the average energy is the same for all the black hole mass cases. However, we can find that the mass of the black hole could affect the speed at which the average energy evolves to its asymptotical value shown in (3.1). The smaller the black hole mass is, the faster the average energy evolves to its asymptotical value.
As discussed for the $\Delta<A$ case above, we can arrive at the relevant conclusion directly from the expression of the average energy of the QB shown in the Eqs. (3.1) and (3.1). When the local Hawking temperature is fixed, the black hole mass can only affect the incoherent relaxation and dephasing rates shown in the Eqs. (3.1) and (3.1), thus can not lead to qualitatively behavior. Note that this result of QB here is quite different from that of the Fisher information when the mass is changed shown in Ref. FIBH , where the behavior of the Fisher information can present qualitatively different behavior when the black hole mass is changed.